# A Comprehensive Survey in LLM(-Agent) Full Stack Safety: Data, Training and Deployment

### v1

@@ 6.4 Agent Safety Evaluation
6.4.2 Module-Specific Benchmarks.
@@ 6.4.5 LLM Deployment Perspective
(1) Attack strategies will become more structured and semantically aligned. (i) Black-box attacks may evolve through agent-based optimization, enabling sentence-level jailbreaks with clearer intent and higher success rates. (ii) To overcome the limitations of token-level gradient attacks, future work may focus on generating semantically consistent adversarial prompts less detectable by perplexity-based defenses. (iii) Open-source models will serve as surrogates for closed models, allowing attackers to replicate decision boundaries before launching white-box attacks. (iv) Variants from fine-tuning pipelines may leak private information through cross-model comparison, introducing version-aware privacy risks.
@@ 6.4.6 Agent Roadmap
Agent The evolution of LLM-based agents originated from role-playing paradigms [853, 747, 854, 855], where researchers investigated organizational structures, role allocation mechanisms, and implementation workflows for task-oriented agents in various social contexts. These systematic explorations not only demonstrated agents’ potential in addressing human societal challenges but also spawned interdisciplinary research programs spanning sociology, organizational theory, and psychology. As the field advanced, research focus shifted toward automated agent workflows [856, 741, 857, 806], domain-specific methods for embodied intelligence, and the development of agent capabilities in tool utilization and memory management. Through this progression, agent systems have emerged as a transformative paradigm for automating human social processes, gaining significant recognition as a viable solution for complex societal automation.
@@ 7 Safety in LLM-Based Application
Hallucination Despite their powerful text generation capabilities, LLMs exhibit hallucination phenomena, generating inaccurate, misleading, or entirely fictitious content [861, 862, 863, 864]. In high-stakes fields such as healthcare, law, and finance, unreliable AI-generated content can directly lead to incorrect decision-making. For instance, an AI medical assistant providing inaccurate diagnoses could mislead doctors or patients [687], while a financial AI model offering flawed market predictions might result in substantial investment losses or even financial crises [865]. Ensuring the accuracy and reliability of AI-generated content is fundamental to the commercialization of LLM-based applications.
@@
Privacy Data privacy concerns [866] represent another significant challenge in LLM deployment [767]. Training these models requires vast amounts of text data, which may include personal information, corporate secrets, and medical records [867]. If an LLM inadvertently leaks sensitive training data or lacks robust access control mechanisms, users’ private information could be exploited or misused. In corporate settings, LLMs may unintentionally expose confidential documents or sensitive customer data, leading to severe compliance and legal risks. Moreover, inference-time attacks [868], such as membership inference and model extraction, can further expose sensitive data by allowing adversaries to infer training set membership or replicate model behavior. Therefore, LLM-based applications must incorporate data protection measures and privacy-preserving techniques like differential privacy and query rate limiting to mitigate information leakage risks.
@@
Robustness Prompt injection [500] and jailbreak [591] risks pose additional security threats. Attackers can craft adversarial prompts to bypass security restrictions, causing the model to generate harmful or unauthorized content. For example, in chatbot systems, malicious users could manipulate LLMs to generate hate speech, disinformation, or even harmful instructions. Similarly, in AI-powered coding assistants such as GitHub Copilot, attackers may exploit LLMs to produce code with security vulnerabilities, potentially serving as backdoors for future cyberattacks. Developing robust security defenses to prevent LLMs from being misused in real-world applications is crucial for AI safety.
@@
Copyright Another pressing concern is intellectual property and copyright protection [869, 870, 871]. LLMs are trained on vast datasets that often include copyrighted texts, source code, and artistic works, raising potential infringement risks. When generating content, LLMs may inadvertently replicate or closely mimic copyrighted material, leading to legal disputes. For instance, AI-powered writing tools might generate articles resembling published works, while coding assistants could produce open-source code snippets without proper licensing [872]. This not only raises concerns about content originality but also introduces legal and ethical dilemmas. Addressing these challenges requires watermarking, provenance tracking, and clear copyright attribution mechanisms to ensure responsible AI-generated content management.
@@
Ethical and Social Responsibility Beyond technical concerns, ethical and social responsibility are also critical factors in large-scale LLM deployment. Due to biases in training data, LLMs may generate content that reinforces stereotypes, gender discrimination, or racial biases [873]. In sectors such as hiring, finance, and healthcare, biased AI-generated recommendations could exacerbate existing inequalities and lead to unfair decision-making. Moreover, as LLMs become increasingly integrated into virtual assistants, social media, and news distribution platforms, concerns over AI-generated misinformation, transparency, and accountability are growing. Building fair, transparent, and trustworthy AI governance frameworks is thus essential to mitigating AI-induced social risks.
@@
Governance As governments worldwide strengthen AI regulations, LLM-related legal and compliance requirements are evolving rapidly. The EU AI Act classifies LLMs as high-risk AI systems, requiring developers to provide transparency reports and risk control mechanisms [874]. China’s Generative AI Regulations mandate AI-generated content to align with ethical standards and undergo governmental scrutiny [875]. In the United States, regulatory discussions emphasize AI transparency and data privacy protections, urging businesses to establish responsible AI practices [876]. These policy developments indicate that LLM-based applications must comply with regional regulations while maintaining a balance between compliance and innovation.
@@ 9 Conclusion
Through an exhaustive review of over 700+ papers, we systematically examined and organized the safety issues spanning key stages of LLM production, deployment, and use, including data generation, alignment techniques, model editing, and LLM-based agent systems. Our findings highlight the critical vulnerabilities at each stage, such as privacy risks, toxic data, harmful fine-tuning attacks, and deployment challenges. The safety of LLMs is a multifaceted issue requiring careful attention to data integrity, model alignment, and post-deployment security measures. Moreover, we propose promising directions for future research, including improvements in data safety, alignment techniques, and defense mechanisms for LLM-based agents. This work is vital for guiding future efforts to make LLMs safer and more reliable, especially as they become increasingly integral to various industries and applications. Ensuring robust security across the entire LLM lifecycle is crucial for their responsible and effective deployment in real-world scenarios.

### v2

@@ 6.5.1 Attack-Specific Benchmarks
6.5.2 Module-Specific Benchmarks
@@ 6.5.5 LLM Deployment Perspective
(1) Attack strategies will become more structured and semantically aligned. (i) Black-box attacks may evolve through agent-based optimization, enabling sentence-level jailbreaks with clearer intent and higher success rates. (ii) To overcome the limitations of token-level gradient attacks, future work may focus on generating semantically consistent adversarial prompts that are less detectable by perplexity-based defenses. (iii) Open-source models will serve as surrogates for closed models, allowing attackers to replicate decision boundaries before launching white-box attacks. (iv) Variants from fine-tuning pipelines may leak private information through cross-model comparison, introducing version-aware privacy risks.
@@ 6.5.6 Agent Roadmap
Agent. The evolution of LLM-based agents originated from role-playing paradigms [931, 794, 932, 933], where researchers investigated organizational structures, role allocation mechanisms, and implementation workflows for task-oriented agents in various social contexts. These systematic explorations not only demonstrated agents’ potential in addressing human societal challenges but also spawned interdisciplinary research programs spanning sociology, organizational theory, and psychology. As the field advanced, research focus shifted toward automated agent workflows [934, 788, 935, 853], domain-specific methods for embodied intelligence, and the development of agent capabilities in tool utilization and memory management. Through this progression, agent systems have emerged as a transformative paradigm for automating human social processes, gaining significant recognition as a viable solution for complex societal automation.
@@ 7 Safety in LLM-Based Application
Truthfulness. Despite their powerful text generation capabilities, LLMs exhibit hallucination phenomena, generating inaccurate, misleading, or entirely fictitious content [939, 940, 941, 942, 943]. Unlike traditional errors, hallucinations are often subtle and linguistically plausible, making them especially dangerous in real-world applications. This challenge is exacerbated in high-stakes domains such as healthcare, law, and finance, where misleading AI-generated information can directly affect human safety and economic stability. For example, an LLM-powered clinical assistant may suggest nonexistent diseases or cite unverified treatments, posing risks to patients [732, 944], while financial advisors powered by LLMs might generate persuasive but flawed market forecasts, leading to significant capital misallocation or systemic financial vulnerabilities [945]. Specifically, hallucination is not merely a surface-level output flaw but a systemic artifact rooted in the model’s training dynamics and the nature of its data. Specifically, hallucination can stem from three compounding factors: (1) semantic overgeneralization due to exposure to noisy, unverified, or synthetic pretraining corpora; (2) objective misalignment, where maximum-likelihood or reinforcement-based training prioritizes coherence and helpfulness over factual accuracy; and (3) latent distribution shifts between pretraining and deployment-time inputs, particularly under long-tail or adversarial queries [946, 947]. These factors jointly reinforce spurious correlations and amplify unsupported generations, even in otherwise well-aligned models. In sum, hallucination represents a critical bottleneck for the reliable deployment of LLMs. Its mitigation is foundational not only for improving user trust but also for enabling the safe integration of LLMs into high-stakes decision-making workflows.
@@
Privacy. Data privacy concerns [948] represent another significant challenge in LLM deployment [814, 949]. Training these models requires vast amounts of text data, which may include personal information, corporate secrets, and medical records [950]. If an LLM inadvertently leaks sensitive training data or lacks robust access control mechanisms, users’ private information could be exploited or misused. In corporate settings, LLMs may unintentionally expose confidential documents or sensitive customer data, leading to severe compliance and legal risks. Moreover, inference-time attacks [951], such as membership inference and model extraction, can further expose sensitive data by allowing adversaries to infer training set membership or replicate model behavior. Therefore, LLM-based applications must incorporate data protection measures and privacy-preserving techniques like differential privacy and query rate limiting to mitigate information leakage risks.
@@
Robustness. Prompt injection [536] and jailbreak [629] risks pose additional security threats. Attackers can craft adversarial prompts to bypass security restrictions, causing the model to generate harmful or unauthorized content. For example, in chatbot systems, malicious users could manipulate LLMs to generate hate speech, disinformation, or even harmful instructions. Similarly, in AI-powered coding assistants such as GitHub Copilot, attackers may exploit LLMs to produce code with security vulnerabilities, potentially serving as backdoors for future cyberattacks. Developing robust security defenses to prevent LLMs from being misused in real-world applications is crucial for AI safety.
@@
Copyright. Another pressing concern is intellectual property and copyright protection [952, 953, 954]. LLMs are trained on vast datasets that often include copyrighted texts, source code, and artistic works, raising potential infringement risks. When generating content, LLMs may inadvertently replicate or closely mimic copyrighted material, leading to legal disputes. For instance, AI-powered writing tools might generate articles resembling published works, while coding assistants could produce open-source code snippets without proper licensing [955]. This not only raises concerns about content originality but also introduces legal and ethical dilemmas. Addressing these challenges requires watermarking [956, 957], provenance tracking, and clear copyright attribution mechanisms to ensure responsible AI-generated content management [178].
@@
Ethical and Social Responsibility. Beyond technical concerns, ethical and social responsibility are also critical factors in large-scale LLM deployment. Due to biases in training data, LLMs may generate content that reinforces stereotypes, gender discrimination, or racial biases [958, 959]. In sectors such as hiring, finance, and healthcare, biased AI-generated recommendations could exacerbate existing inequalities and lead to unfair decision-making. Moreover, as LLMs become increasingly integrated into virtual assistants, social media, and news distribution platforms, concerns over AI-generated misinformation, transparency, and accountability are growing. Building fair, transparent, and trustworthy AI governance frameworks is thus essential to mitigating AI-induced social risks.
@@
Governance. As governments worldwide strengthen AI regulations, LLM-related legal and compliance requirements are evolving rapidly. The EU AI Act classifies LLMs as high-risk AI systems, requiring developers to provide transparency reports and risk control mechanisms [960]. China’s Generative AI Regulations mandate AI-generated content to align with ethical standards and undergo governmental scrutiny [961]. In the United States, regulatory discussions emphasize AI transparency and data privacy protections, urging businesses to establish responsible AI practices [962]. These policy developments indicate that LLM-based applications must comply with regional regulations while maintaining a balance between compliance and innovation.
@@ 9 Conclusion
Through an exhaustive review of over 900+ papers, we systematically examined and organized the safety issues spanning key stages of LLM production, deployment, and use, including data generation, alignment techniques, model editing, and LLM-based agent systems and LLM-based applications. Our findings highlight the critical vulnerabilities at each stage, such as privacy risks, toxic data, harmful fine-tuning attacks, and deployment challenges. The safety of LLMs is a multifaceted issue requiring careful attention to data integrity, model alignment, and post-deployment security measures. Moreover, we propose promising directions for future research, including improvements in data safety, alignment techniques, and defense mechanisms for LLM-based agents. This work is vital for guiding future efforts to make LLMs safer and more reliable, especially as they become increasingly integral to various industries and applications. Ensuring robust security across the entire LLM lifecycle is crucial for their responsible and effective deployment in real-world scenarios.